OpenStudy (anonymous):

G = xw over y + z solve for x i just need help with the steps, if you could please help ?

OpenStudy (anonymous):

@iGreen ?

OpenStudy (anonymous):

\[G= \frac{ xw }{ y+z}\] is this what it looks like?

OpenStudy (anonymous):

yes, it is!

OpenStudy (anonymous):

ok, so you would multiply both sides by y+z and get g(y+z)=xw then you can distribute the g if you want. after that you can divide both sides by w to get x alone \[\frac{ g(y+z) }{ w }=x\]

OpenStudy (anonymous):

thank you so much, your a life saver!!

OpenStudy (anonymous):

No problem! Glad i could help!!